Problem 1: While you were writing this quiz, an announcement of an important scientific discovery was made: a German chemist Hanz Richerberg has discovered a new inexpensive method of turning iron into gold. What will be the effect of this announcement on the basis of April futures gold contracts?

  1. The bases will significantly increase
  2. The basis will significantly decrease
  3. The effect on the basis will be small

Ans: C

Problem 2: Assume that the expected short-term interest rate in the future is equal to the current short-term interest rate. Then, according to the Liquidity Preference Theory,

  1. Forward short-term interest rates are equal to the current long-term interest rates
  2. Forward short-term interest rates are lower than the current long-term interest rates
  3. Forward short-term interest rates are higher than the current long-term interest rates
  4. No prediction can be made about the relation between forward short-term interest rates and current long-term interest rates

Ans: C

Question 3: An investor owns 1,000 shares of AAA inc. currently worth $35 per share. Assume both call and put options on AAA inc. are traded on the market.  How the investor can hedge his portfolio using these options?

  1. Buy 1000 call options only
  2. Sell 1000 call options only
  3. Buy 1000 put options only
  4. Sell 1000 put options only
  5. Either buy 1000 call options or sell 1000 put options
  6. Either buy 1000 put options or sell 1000 call options

Ans: C

Question 4: The initial margin for a future contract on 10,000 bushels of corn is $4,000 and the maintenance margin for the same contract is $3,500. At the beginning of the day you had $4,250 on your margin account. During the trading day the futures price went down by $0.08 while the spot price went down only by $0.06. If you have one long position in futures contracts, which of the following is true:

  1. You will receive a margin call and you will need to add $800 to your margin account
  2. You will receive a margin call and you will need to add $600 to your margin account
  3. You will receive a margin call and you will need to add $550 to your margin account
  4. You will receive a margin call and you will need to add $350 to your margin account
  5. You will receive a margin call and you will need to add $50 to your margin account
  6. You will not receive a margin call

Ans: C

Please explain each of the answers
Answer #1

1. Basis risk occurs when price of the underlying and the hedge (Gold futures in this case) move in opposite directions thus reducing the effectiveness of hedge. The given scientific discovery may increase the supply of gold in the market and the prices of both spot and futures of gold will move in one direction. Therefore will not affect the basis significantly.

2. As per liquidity preference theory, money in hand now is more valuable than money in future.

When $100 is invested, say, after 5 year at future short term interest rate for a specific tenor, the proceeds earned on maturity when discounted (at long term interest rate) to the present will be lower in value than when the same $100 is invested today for the same tenor. Therefore in order to compensate for the value lost, short term forward rate must be higher than the long term interest rate.

3. Investor owns 1000 shares at $35. He will have a loss if prices fall below $35. (We are assuming zero option prices).

Let us assume share price drops to $30. If investor buy a put option he will have right to execute the option and sell the share at $35 (i.e. at higher price then the market share price ). Thus total return from stock + option will be=(s1-s0)+(s1-x)= (30-35)+(35-30)=0

S0=35 s1=30 option stike price, x=35

Thus position is hedged

Again if prices are above $35, let's say $40, there will be total gain of $5 in spot and option will not be executed

4. Initial margin= 4000 Maintenance margin= 3500

As future prices are gone down by 0.08 per future, total end of day loss is 10,000*0.08= $800

Current Balance= $4250

After mark to market Marin level= $4250-$800 = $3450

If balance goes below maintenance margin, margin call is made to replenish the initial margin. Thus call of $550 will be made